OpenStudy (anonymous):

An Astronomical Unit is defined as 9.3e7 miles or the distance of the Earth from the Sun when Mars is closest to Earth it is 1.3814 AU. Assuming we can get a spacecraft to travel at the same velocity as Voyager 1(17km/sec) how many years years would it take a spacecraft full of people that want to live on Mars ro fet from Earth to Mars. Please show the process.

OpenStudy (noelgreco):

d=rt

OpenStudy (anonymous):

1 AU = 9.3 x 10^7 miles 1.3814 * (1 AU) = 1.3814 * (9.3 x 10^7) miles 1.3814 AU = 1.284702 x 10^8 miles 1 mile = 1.609 kilometers 1.609 * (1.284702 x 10^8 miles) = 2.06708552 x 10^8 kilometers 2.06708552 x 10^8 kilometers * (1 second / 17 kilometers) = 1.2159327 x 10^7 seconds 1 year = 3.1536 x 10^7 seconds 1.2159327 x 10^7 seconds * (1 year / 3.1536 x 10^7 seconds) = 0.385569717 years So about 140.82 days.